# Small maximal clusters are very unlikely in critical random graphs

### v1

@@ 1.3 Some useful tools
1.4 Notation and structure of the paper
Notation.
We let =$\mathbb{N}=\{1,2,\dots,\}$ and define ${\mathbb{N}}_{0}\coloneqq{\mathbb{N}}\cup\{0\}$. For $n\in\mathbb{N}$ we set $[n]\coloneqq\{1,2,\ldots,n\}$. Given two sequences of real numbers $(x_{n})_{n\geq 1}$ and $(y_{n})_{n\geq 1}$ we write: $(1)$ $x_{n}=O(y_{n})$ if there exist $N\in\mathbb{N}$ and $C\in[0,\infty)$ such that $|x_{n}|\leq C|y_{n}|$ for all $n\geq N$; (2) either $x_{n}=o(y_{n})$ or $x_{n}\ll y_{n}$ if $x_{n}/y_{n}\rightarrow 0$ as $n\rightarrow\infty$; and (3) either $x_{n}=\Theta(y_{n})$ or $x_{n}\asymp y_{n}$ if $x_{n}=O(y_{n})$ and $y_{n}=O(x_{n})$. We often write $c,C$ to denote constants in $(0,\infty)$, and use these letters many times in a single proof even though their actual values may change from line to line.
When talking about random variables, the abbreviation iid means independent and identically distributed. Moreover, given random variables $X$ and $Y$, we write $\leq_{sd}$ if $\mathbb{P}(X\geq z)\leq\mathbb{P}(Y\geq z)$ for every $z\in\mathbb{R}$, and say that $X$ is stochastically dominated by $Y$. We write $\text{Bin}(N,q)$ to denote the binomial distribution of parameters $N\in\mathbb{N}_{0},q\in[0,1]$ and $\text{Unif}([0,1])$ to denote the uniform distribution on $[0,1]$. Given a distribution function $F$, we use the notation $X=_{d}F$ to mean that $X$ has the distribution $F$.
Let $\mathbb{G}=(V,E)$ be any (undirected, possibly random) multigraph. Given two vertices $u,v\in V$, we write $u\sim v$ if $\{u,v\}\in E$ and say that vertices $u$ and $v$ are neighbours. We often write $uv$ as shorthand for the edge $\{u,v\}$. We write $u\leftrightarrow v$ if there exists a path of occupied edges connecting vertices $u$ and $v$, where we adopt the convention that $v\leftrightarrow v$ for every $v\in V$. We denote by $\mathcal{C}(v)\coloneqq\{u\in V:u\leftrightarrow v\}$ the component containing vertex $v\in V$. We define a largest component $\mathcal{C}_{\max}(\mathbb{G})$ of $\mathbb{G}$ to be some cluster $\mathcal{C}(v)$ for which $|\mathcal{C}(v)|$ is maximal, so that $|\mathcal{C}_{\max}(\mathbb{G})|=\max_{v\in V}|\mathcal{C}(v)|$.
Structure of the paper.
The rest of the paper is organised as follows. In Section 2 we prove Theorem 1.1, whereas in Section 3 we describe how to modify the martingale argument of [28, 27] with the purpose of obtaining stretched exponential estimates in several critical models.
@@ 1.4 Notation and structure of the paper
Here the goal is to show that $|\mathcal{C}_{\max}(\mathbb{G})|\geq n^{2/3}/A$ with probability at least of order $1-e^{-A^{3/2}}$ when either $\mathbb{G}=\mathbb{G}(n,p)$ and $p=(1+\lambda n^{-1/3})/n$ or $\mathbb{G}=\mathbb{G}(n,d,p)$ and $p=(1+\lambda n^{-1/3})/(d-1)$ (with $\lambda\in\mathbb{R},d\geq 3$ fixed).
We begin by providing an heuristic derivation of these estimates and we do so in terms of the $\mathbb{G}(n,p)$ random graph; the proof for the $\mathbb{G}(n,d,p)$ model follows the same line of thoughts.
@@
In order to study $|\mathcal{C}_{\max}(\mathbb{G}(n,p))|$ we use (as it is standard, see e.g. [20]) an exploration process, which is an algorithmic procedure to sequentially reveal the connected components of $\mathbb{G}(n,p)$ and that reduces the study of component sizes to the analysis of the trajectory of a (self-interacting) discrete-time, non-negative stochastic process.
Roughly speaking, an exploration process (in this setting) works as follows. We start by specifying an ordering of the vertices. Then, at time $t=0$, one of the nodes, say $u$, is declared active, whereas all the remaining nodes are unseen and there are no explored vertices. At time $t=1$, we reveal the unseen neighbours of $u$, which are declared active. The step terminates by declaring $u$ explored. Then we iterate the procedure: at time $t=2$, we select one of the active nodes (if any), say $v$, and reveal its unseen neighbours, which are declared active, whereas $v$ itself becomes explored. If at the start of some step $t$ the set of active vertices is empty, then we pick the first (with respect to the ordering fixed at the very beginning) unseen vertex (if any), say $w$, reveal its unseen neighbours, which are declared active and change the status of $w$ to explored. The procedure terminates when all the vertices are in status explored.
We notice that the exploration of a connected component, started at some time $t$, continues as long as there are active vertices, say until time $t+k$, so that the size of the underlying cluster equals $k$, the number of steps of the procedure (starting from time $t$) for which the set of active vertices is non-empty.
Denote by $Y_{t}$ the number of active vertices at the end of step $t$ in the exploration process. Following the above description, we use such a procedure to bound (from above) the probability that $\mathcal{C}_{\max}(\mathbb{G}(n,p))$ contains less than $n^{2/3}/A$ vertices by the probability that the positive excursions of $Y_{t}$ never last for more than $n^{2/3}/A$ steps.
More precisely, defining $t_{i}$ to be the first time $t>t_{i-1}$ at which $Y_{t}=0$ (prior to the end of the procedure) and denoting by $\mathcal{C}_{i}$ the $i$-th explored cluster, whose exploration starts at time $t_{i}$ so that $|\mathcal{C}_{i}|=t_{i}-t_{i-1}$, we can bound from above the probability that $\mathcal{C}_{\max}(\mathbb{G}(n,p))$ contains less than $n^{2/3}/A$ vertices by the probability that each excursion length $t_{i}-t_{i-1}$ is at most $n^{2/3}/A$:
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G}(n,p))|<n^{2/3}/A)=\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall i\in[N]),$ (11)
@@
$\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall i\in[N])\leq\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall L/2\leq i\leq L)$
and
$\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall L/2\leq i\leq L)=\prod_{i=L/2}^{L}\Big{(}1-\mathbb{P}(t_{i}-t_{i-1}\geq T|t_{j}-t_{j-1}<T\text{ }\forall L/2\leq j\leq i-1)\Big{)}.$ (12)
We claim that, for each $L/2\leq i\leq L$,
@@
Observe that, if (13) were true, then (recalling the value of $T$ and using the classical inequality $1+x\leq e^{x}$) the expression on the right-hand side of (12) would satisfy
$\prod_{i=L/2}^{L}\Big{(}1-\mathbb{P}(t_{i}-t_{i-1}\geq T|t_{j}-t_{j-1}<T\text{}\forall L/2\leq j\leq i-1)\Big{)}\\ \lesssim\Big{(}1-T^{-1/2}\Big{)}^{L/2}\leq\exp(-L/(2T^{1/2}))\lesssim\exp(-A^{3/2}),$
which is the desired bound. Hence we are left with the problem of showing (13). To this end, recall the description of the exploration process given earlier. It is clear that, if at time $t-1$ there is at least one active node then, denoting by $\eta_{t}$ the number of unseen vertices which become active at step $t$, we have $Y_{t}=Y_{t-1}+\eta_{t}-1$: indeed, the number of active vertices at time $t$ equals the number of active nodes at time $t-1$, plus the newly discovered (necessarily unseen) neighbours of the vertex under investigation at step $t$, minus one (since the status of the vertex under exam passes from active to explored). Consequently, if the number of active vertices stays positive from a time $s$ until a time $s+t$, then we can write $Y_{s+t}=1+\sum_{k=1}^{t}(\eta_{s+k}-1)$. Moreover, conditional on the history of the exploration process until time $s+k-1$, it is clear from the model definition and the description of the exploration process that the random variable $\eta_{s+k}$ has the $\text{Bin}(n-s-k+1-Y_{s+k-1},p)$ distribution (where $n-s-k+1-Y_{s+k-1}$ is the number of unseen vertices at the start of step $t$).
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|t_{j}-t_{j-1}<T\text{ }\forall L/2\leq j\leq i-1)\approx\mathbb{P}\Big{(}\sum_{k=1}^{t}(\text{Bin}_{k}(n-t_{i-1}-k,p)-1)>0\text{ }\forall t\leq T\Big{)}.$
Let us assume that $t_{L}\lesssim LT^{1/2}$; then, for each $L/2\leq i\leq L$, the exploration of the $i$-th cluster $\mathcal{C}_{i}$ starts before time $LT^{1/2}$, which implies that, when starting the exploration of $\mathcal{C}_{i}$, the number of explored vertices is $\lesssim LT^{1/2}$ (for each $L/2\leq i\leq L$). Under this assumption, after noticing that $k\leq T=n^{2/3}/A\ll A^{1/2}n^{2/3}=LT^{1/2}$ (whence $LT^{1/2}+k\leq 2LT^{1/2}$ for all large enough $n$), we can couple the $\text{Bin}_{k}(n-t_{i-1}-k,p)$ random variables appearing within the above probability with independent $\text{Bin}_{k}(n-2LT^{1/2},p)$ random variables in such a way that
$\text{Bin}_{k}(n-t_{i-1}-k,p)\geq\text{Bin}_{k}(n-2LT^{1/2},p)\text{ for each }k\in[T].$
@@
$\mathbb{P}\Big{(}\sum_{k=1}^{t}(\text{Bin}_{k}(n-t_{i-1}-k,p)-1)>0\text{ }\forall t\leq T\Big{)}\gtrsim\mathbb{P}\Big{(}\sum_{k=1}^{t}(\text{Bin}_{k}(n-2LT^{1/2},p)-1)>0\text{ }\forall t\leq T\Big{)}.$ (14)
Now the process considered in the last probability is a random walk with a small negative drift. Such a small drift does not affect the behaviour of the process, which from a practical point of view can be regarded as a $\mathbb{Z}$-valued, mean-zero random walk. More specifically, we can approximate the probability in (14) by $\mathbb{P}(S_{t}>0\text{ }\forall t\leq T)$, where $S_{t}$ a random walk with iid $\text{Bin}(n,p)-1$ increments. It is well-known that a mean-zero random walk with iid increments having a finite second moment stays above zero for $T$ steps with probability $\asymp T^{-1/2}$, thus establishing the claim (13).
@@
We let $t_{0}\coloneqq 0$ and recursively we define $t_{i}$ to be the first time after $t_{i-1}$ at which the number of active vertices hits zero; that is,
@@
where we denote by $N$ the (random) number of excursions (i.e. the number of clusters). Setting $T\coloneqq\lceil n^{2/3}/A\rceil$, we write
@@
where $L\leq j\leq 2L$ and $B$ is some large positive constant (independent of $A,n,\lambda$). The event $\mathcal{G}_{j}$ tells that we start exploring the $j$-th cluster before time $BLT^{1/2}$, which also means that when starting exploring the $j$-th cluster, the number of explored vertices is at most $BLT^{1/2}$ (recall that at each step of the exploration exactly one vertex becomes explored). A good control on the $t_{j}$’s is needed when bounding from below the probability that $t_{i}-t_{i-1}$ is larger than $T$, which is the same as the probability that the number of active vertices stays above zero for $T$ steps starting from time $t_{i-1}$. If the number of explored vertices were ‘too large’ when starting the exploration of $\mathcal{C}_{i}$, then the latter probability would be ‘too small’, because the negative drift of the increments $\eta_{t_{i-1}+k}-1$ would be ‘too large’ (in absolute value) to keep the process above zero for $T$ steps.
@@
$\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L)\leq\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L})+\mathbb{P}(\mathcal{G}^{c}_{2L}).$
The next lemma controls the probability of the (bad) event $\mathcal{G}^{c}_{2L}$.
Lemma 2.1.
@@
($\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L}).$ (17)
Let us define $\mathcal{F}_{j}\coloneqq\mathcal{F}_{t_{j}}$ for $j\in[N]$, the $\sigma$-algebra consisting of all the information collected by the exploration process by time $t_{j}$, i.e. by the time at which we start the exploration of the $j$-th component.
With the next result we show that, on the event $\mathcal{G}_{i-1}$, we have $t_{i}-t_{i-1}\geq T$ with probability $\gtrsim T^{-1/2}$ for each $L+1\leq i\leq 2L$.
Lemma 2.2.
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{i-1})\geq\frac{c_{0}}{T^{1/2}}$
@@
It is now very easy to show that the expression in (17) is at most $\exp(-cA^{3/2})$ (for some $c>0$). Indeed, using the tower property of conditional expectation, together with the fact that $\mathcal{G}_{j}\subset\mathcal{G}_{j-1}$, we can write
($\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L})\\ \leq\mathbb{E}\Big{[}\mathbbm{1}_{\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L-1\}\cap\mathcal{G}_{2L-1}}\big{(}1-\mathbb{P}(t_{2L}-t_{2L-1}\geq T|\mathcal{F}_{2L-1})\big{)}\Big{]}.$ (18)
@@
Iterating and using the classical inequality $1+x\leq e^{x}$ (which is valid for all $x\in\mathbb{R}$), we conclude that for all large enough $A$ and $n$
@@
This concludes the proof of Theorem 1.1 for the $\mathbb{G}(n,p)$ random graph. What is left to do is to prove Lemmas 2.1 and 2.2; this is done in the following section.
2.2.1 Proofs of Lemmas 2.1 and 2.2
@@
Proof of Lemma 2.1.
Recall that $\mathcal{G}_{2L}\coloneqq\{t_{2L}\leq BLT^{1/2}\}$. To bound from above the probability of the complementary event, we proceed as follows. Let $S_{0}\coloneqq 1$ and define recursively $S_{t}\coloneqq S_{t-1}+\eta_{t}-1$, where the random variable $\eta_{i}$, conditional on the history of the exploration process until time $i-1$ (denoted here by $\mathcal{H}_{i-1}$), has the $\text{Bin}(n-(i-1)-Y_{i-1},p)$ distribution.
The $\mathbb{Z}$-valued random process $(S_{t})_{t\in\mathbb{N}_{0}}$ is closely related to the process of active vertices $(Y_{t})_{t\in\mathbb{N}_{0}}$; in particular, they coincide up to time $t_{1}$ (the time at which we completes the exploration of the first cluster). However, at time $t_{1}+1$, we have $Y_{t_{1}+1}=\eta_{t_{1}+1}$ whereas $S_{t_{1}+1}=\eta_{t_{1}+1}-1$. The two processes $Y_{t_{1}+t}$ and $S_{t_{1}+t}$ have the same increments for $t_{1}+1<t\leq t_{2}$, but then again $Y_{t_{2}+1}=\eta_{t_{2}+1}$ whereas $S_{t_{2}+1}=\eta_{t_{2}+1}-2$, and so on
Then $(-\min_{j\in[t]}S_{j})+1$ denotes the number of disjoint clusters that have been fully explored by time $t\in[n]$ (see e.g. Chapter 5.2.4 in [20]). When $Y_{t}=0$, we have fully explored a cluster, and this occurs precisely when $S_{t}=\min_{j\in[t-1]}S_{j}-1$, i.e., the process $(S_{t})_{t\in[n]}$ reaches a new all-time minimum.
Now observe that, if $t_{2L}>BLT^{1/2}$, then the number of components that are fully explored up to time $\lfloor BLT^{1/2}\rfloor$ is smaller than $2L$ (exactly because $t_{2L}$ is the time at which we start exploring the $(2L)$-th cluster). Therefore we can bound
$\mathbb{P}(\mathcal{G}^{c}_{2L})\leq\mathbb{P}\Big{(}-\min_{j\leq\lfloor BLT^{1/2}\rfloor}S_{j}<2L-1\Big{)}=\mathbb{P}\Big{(}\max_{j\leq\lfloor BLT^{1/2}\rfloor}(-S_{j})<2L-1\Big{)}\leq\mathbb{P}\big{(}S_{\lfloor BLT^{1/2}\rfloor}>-2L\big{)}.$ (20)
@@
$S_{\lfloor BLT^{1/2}\rfloor}=1+\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}(\eta_{i}-1)\text{ and }(\eta_{i}|\mathcal{H}_{i-1})=_{d}\text{Bin}(n-(i-1)-Y_{i-1},p)\leq_{sd}\text{Bin}(n-(i-1),p).$
@@
$\mathbb{P}(S_{\lfloor BLT^{1/2}\rfloor}>-2L)\leq\mathbb{P}\Big{(}\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}(\delta_{i}-1)>-2L\Big{)}.$
@@
where $c=c(B)>0$ is a constant which depends on $B$. Since $TP=O(A^{-1}n^{-1/3})$,
@@
Therefore, going back to (21) and setting
$\ell\coloneqq\lfloor BLT^{1/2}\rfloor(1-np)+(Mp)/2-3L,$
we can bound
@@
A simple computation shows that, for all large enough $A$ and $n$,
$(Mp)/2-4L\geq\frac{B^{2}L^{2}T}{2n}-5L=L\Big{(}\frac{B^{2}LT}{2n}-5\Big{)}\geq L(B^{2}/2-5-o(1))\geq L(B^{2}/2-6),$

### v3

@@ 1.4 Some useful tools
Here the goal is to show that $|\mathcal{C}_{\max}(\mathbb{G})|\geq n^{2/3}/A$ with probability at least $1-\exp(-\Omega(A^{3/2}))$ when either $\mathbb{G}=\mathbb{G}(n,p)$ and $p=(1+\lambda n^{-1/3})/n$, or $\mathbb{G}=\mathbb{G}(n,d,p)$ and $p=(1+\lambda n^{-1/3})/(d-1)$, with $3\leq d\in\mathbb{N}$ fixed. We begin by providing an heuristic derivation of these estimates and we do so in terms of the $\mathbb{G}(n,p)$ random graph; the proof for the $\mathbb{G}(n,d,p)$ model is similar.
@@
In order to study the size of $\mathcal{C}_{\max}(\mathbb{G}(n,p))$ we use (as it is standard, see e.g. [20]) an exploration process, which is an algorithmic procedure to sequentially reveal the connected components of a simple, undirected graph, that we apply to $\mathbb{G}(n,p)$ and which reduces the study of component sizes to the analysis of the trajectory of a (self-interacting) discrete-time, non-negative stochastic process.
Roughly speaking, the exploration process (in this setting) works as follows. We start by specifying an ordering of the vertices. Then, at time $t=0$, one of the nodes, say $u$, is declared active, whereas all the remaining nodes are unseen and there are no explored vertices. At time $t=1$, we reveal the unseen neighbours of $u$, which are declared active. The step terminates by declaring $u$ explored. Then we iterate the procedure: at time $t=2$, we select one of the active nodes (if any), say $v$, and reveal its unseen neighbours, which are declared active, whereas $v$ itself becomes explored. If at the start of some step $t$ the set of active vertices is empty, then we pick the first (with respect to the ordering fixed at the very beginning) unseen vertex (if any), say $w$, reveal its unseen neighbours, which are declared active and change the status of $w$ to explored. The procedure terminates when all the vertices are in status explored.
We notice that the exploration of a connected component, started at some time $t$, continues as long as there are active vertices, say until time $t+k$, so that the size of the underlying cluster equals $k$, the number of steps of the procedure (starting from time $t$) at which the set of active vertices is non-empty. Following the above description, we use such a procedure to bound (from above) the probability that $\mathcal{C}_{\max}(\mathbb{G}(n,p))$ contains less than $n^{2/3}/A$ vertices by the probability that the positive excursions of $Y_{t}$ never last for more than $n^{2/3}/A$ steps, where we write $Y_{t}$ for the number of active vertices at the end of step $t$ in the exploration process.
More precisely, setting $t_{0}\coloneqq 0$ and defining $t_{i}$ to be the first time $t>t_{i-1}$ at which $Y_{t}=0$ (prior to the end of the procedure) and denoting by $\mathcal{C}_{i}$ the $i$-th explored cluster, whose exploration starts at time $t_{i-1}+1$, we see that $|\mathcal{C}_{i}|=t_{i}-t_{i-1}$, we can bound from above the probability that $\mathcal{C}_{\max}(\mathbb{G}(n,p))$ contains less than $n^{2/3}/A$ vertices by the probability that each excursion length $t_{i}-t_{i-1}$ is at most $n^{2/3}/A$:
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G}(n,p))|<n^{2/3}/A)=\mathbb{P}(t_{i}-t_{i-1}<n^{2/3}/A\text{ }\forall i\in[N]),$ (11)
@@
$\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall i\in[N])\leq\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall L/2\leq i\leq L).$
Moreover,
$\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall L/2\leq i\leq L)\leq\prod_{i=L/2+1}^{L}\Big{(}1-\mathbb{P}(t_{i}-t_{i-1}\geq T|t_{j}-t_{j-1}<T\text{ }\forall L/2\leq j\leq i-1)\Big{)}.$ (12)
We claim that, for each $L/2<i\leq L$,
@@
Observe that, if (13) were true, then (recalling the value of $T$ and using the classical inequality $1+x\leq e^{x}$, valid for every real $x$) the expression on the right-hand side of (12) would satisfy
$\prod_{i=L/2+1}^{L}\Big{(}1-\mathbb{P}(t_{i}-t_{i-1}\geq T|t_{j}-t_{j-1}<T\text{ }\forall L/2\leq j\leq i-1)\Big{)}\\ \lesssim\Big{(}1-T^{-1/2}\Big{)}^{L/2}\leq\exp(-L/(2T^{1/2}))=\exp(-\Omega(A^{3/2})),$
which is the desired upper bound. Hence we are left with the problem of showing (13). To this end, recall the description of the exploration process given earlier. It is clear that, if at time $t-1$ there is at least one active node then, letting $u$ denoting one of them and writing $\eta_{t}$ for the number of unseen neighbors of $u$ which become active at step $t$, we have $Y_{t}=Y_{t-1}+\eta_{t}-1$: indeed, the number of active vertices at time $t$ equals the number of active nodes at time $t-1$, plus the newly discovered (necessarily unseen) neighbors of $u$, minus one (since the status of $u$ passes from active to explored). Consequently, if the number of active vertices stays positive from time $s$ until a time $s+t$, then we can write $Y_{s+t}=1+\sum_{k=1}^{t}(\eta_{s+k}-1)$. Moreover, conditional on the history of the exploration process until time $s+k-1$, it is clear from the model definition and the description of the exploration process that the random variable $\eta_{s+k}$ has the $\text{Bin}(n-s-k+1-Y_{s+k-1},p)$ distribution (where $n-s-k+1-Y_{s+k-1}$ is the number of unseen vertices at the start of step $s+k$).
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|t_{j}-t_{j-1}<T\text{ }\forall L/2\leq j\leq i-1)\approx\mathbb{P}\Big{(}\sum_{k=1}^{t}(\text{Bin}_{k}(n-t_{i-1}-k,p)-1)>0\text{ }\forall t<T\Big{)}.$
Let us assume that the exploration of the $(L-1)$-th cluster finishes before time $LT^{1/2}$; that is, $t_{L-1}\lesssim LT^{1/2}$. Under this assumption, after noticing that each $k$ in the above sum of binomials satisfies $k\leq T=n^{2/3}/A\ll A^{1/2}n^{2/3}=LT^{1/2}$ (whence $LT^{1/2}+k\leq 2LT^{1/2}$ for all large enough $n$), we can couple the $\text{Bin}_{k}(n-t_{i-1}-k,p)$ random variables appearing within the above probability with independent $\text{Bin}_{k}(n-2LT^{1/2},p)$ random variables in such a way that
$\text{Bin}_{k}(n-t_{i-1}-k,p)\geq\text{Bin}_{k}(n-2LT^{1/2},p)\text{ for each }k<T.$
@@
$\mathbb{P}\Big{(}\sum_{k=1}^{t}(\text{Bin}_{k}(n-t_{i-1}-k,p)-1)>0\text{ }\forall t<T\Big{)}\gtrsim\mathbb{P}\Big{(}\sum_{k=1}^{t}(\text{Bin}_{k}(n-2LT^{1/2},p)-1)>0\text{ }\forall t<T\Big{)}.$ (14)
Now the process considered in the last probability is a random walk with a small negative drift, as $\mathbb{E}[\text{Bin}_{k}(n-2LT^{1/2},p)]\approx 1-A^{1/2}n^{-1/3}$. Such a small drift does not alter too much the behavior of the process, which from a practical point of view can be regarded as a $\mathbb{Z}$-valued, mean-zero random walk with finite variance. It is well-known that a random walk with these properties stays above zero for $T$ steps with probability $\Theta(T^{-1/2})$, thus establishing the claim (13).
@@
We set $t_{0}\coloneqq 0$ and recursively we define $t_{i}$ to be the first time after $t_{i-1}$ at which the number of active vertices hits zero; that is,
@@
where we denote by $N$ the (random) number of excursions of the process $Y_{t}$ (which is the same as the number of clusters in the random graph). Setting $T\coloneqq\lceil n^{2/3}/A\rceil$, we write
@@
where $L\leq j\leq 2L-1$ and $B$ is some large positive constant (independent of $A,n,\lambda$). The event $\mathcal{G}_{j}$ says that we start exploring the $(j+1)$-th cluster before time $BLT^{1/2}$, which also means that when starting exploring the $(j+1)$-th cluster, the number of explored vertices is at most $BLT^{1/2}$ (recall that at each step of the exploration exactly one vertex becomes explored). A good control on the $t_{j}$’s is needed when bounding from below the probability that $|\mathcal{C}_{i}|=t_{i}-t_{i-1}$ is larger than $T$, which is the same as the probability that the number of active vertices stays above zero for $T$ steps starting from time $t_{i-1}$. Indeed, if the number of explored vertices were ‘too large’ when starting the exploration of $\mathcal{C}_{i}$, then the latter probability (that is, the probability that $Y_{t}>0$ for $T$ steps) would be ‘too small’, because the negative drift of the increments $\eta_{t_{i-1}+k}-1$ would be ‘too large’ (in absolute value) to keep the process above zero for $T$ steps.
@@
$\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L)\leq\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L-1})+\mathbb{P}(\mathcal{G}^{c}_{2L-1}).$
The next result controls the probability of the (bad) event $\mathcal{G}^{c}_{2L-1}$ and shows that it is at most $\exp(-\Omega(A^{3/2}))$.
Proposition 2.1.
@@
($\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L-1}).$ (17)
Let us denote by $\mathcal{F}_{j}$ the $\sigma$-algebra consisting of all the information collected by the exploration process by time $j$; then $\mathcal{F}_{t_{j}}$ represents the information collected by the time at which we finish the exploration of the $j$-th component.
With the next result we show that, on the event $\mathcal{G}_{i-1}$, we have $t_{i}-t_{i-1}\geq T$ with probability at least of order $T^{-1/2}$ for each $L+1\leq i\leq 2L-1$.
Proposition 2.2.
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{t_{i-1}})\geq\frac{c_{0}}{T^{1/2}}$
@@
It is now very easy to show that the expression in (17) is at most $\exp(-cA^{3/2})$ (for some $c>0$). Indeed, using the tower property of conditional expectation, we can write
($\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L-1})\\ =\mathbb{E}\Big{[}\mathbbm{1}_{\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L-1\}\cap\mathcal{G}_{2L-1}}\big{(}1-\mathbb{P}(t_{2L}-t_{2L-1}\geq T|\mathcal{F}_{t_{2L-1}})\big{)}\Big{]}.$ (18)
@@
Iterating and using the classical inequality $1+x\leq e^{x}$ (which is valid for all $x\in\mathbb{R}$) together with the fact that $\mathcal{G}_{j}\subset\mathcal{G}_{j-1}$ for each $j$, we conclude that for all large enough $A$ and $n$
@@
This concludes the proof of Theorem 1.1 for the $\mathbb{G}(n,p)$ random graph. What is left to do is to prove Propositions 2.1 and 2.2; this is done in the following section.
2.2.1 Proofs of Propositions 2.1 and 2.2
@@
Proof of Proposition 2.1.
Recall that $\mathcal{G}_{2L-1}\coloneqq\{t_{2L-1}\leq BLT^{1/2}\}$. To bound from above the probability of the complementary event, we proceed as follows. Let $S_{0}\coloneqq 1$ and define recursively $S_{t}\coloneqq S_{t-1}+\eta_{t}-1$, where the random variable $\eta_{i}$, conditional on the history of the exploration process until time $i-1$ (which we recall is denoted by $\mathcal{F}_{i-1}$), has the $\text{Bin}(n-(i-1)-Y_{i-1},p)$ distribution.
The $\mathbb{Z}$-valued random process $(S_{t})_{t\in\mathbb{N}_{0}}$ is closely related to the process of active vertices $(Y_{t})_{t\in\mathbb{N}_{0}}$; in particular, they coincide up to time $t_{1}$ (the time at which we completes the exploration of the first cluster). However, at time $t_{1}+1$, we have $Y_{t_{1}+1}=\eta_{t_{1}+1}$ whereas $S_{t_{1}+1}=\eta_{t_{1}+1}-1$. The two processes $Y_{t_{1}+t}$ and $S_{t_{1}+t}$ have the same increments for $t_{1}+1<t\leq t_{2}$, but then again $Y_{t_{2}+1}=\eta_{t_{2}+1}$ whereas $S_{t_{2}+1}=\eta_{t_{2}+1}-2$, and so on. Then $(-\min_{j\in[t]}S_{j})+1$ denotes the number of disjoint clusters that have been fully explored by time $t\in[n]$ (see e.g. Chapter 5.2.4 in [20]). When $Y_{t}=0$, we have fully explored a cluster, and this occurs precisely when $S_{t}=\min_{j\in[t-1]}S_{j}-1$, i.e., the process $(S_{t})_{t\in[n]}$ reaches a new all-time minimum.
Now observe that, if $t_{2L-1}>BLT^{1/2}$, then the number of components that are fully explored up to time $\lfloor BLT^{1/2}\rfloor$ is smaller than $12L-$ 1 (precisely because $t_{2L-1}$ is the time at which we finish exploring the $(2L-1)$-th cluster). Therefore we can bound
$\mathbb{P}(\mathcal{G}^{c}_{2L-1})\leq\mathbb{P}\Big{(}-\min_{j\leq\lfloor BLT^{1/2}\rfloor}S_{j}<2L-1\Big{)}=\mathbb{P}\Big{(}\max_{j\leq\lfloor BLT^{1/2}\rfloor}(-S_{j})<2L-1\Big{)}\leq\mathbb{P}\big{(}S_{\lfloor BLT^{1/2}\rfloor}>-2L+1\big{)}.$ (20)
@@
$S_{\lfloor BLT^{1/2}\rfloor}=1+\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}(\eta_{i}-1)\text{ and }(\eta_{i}|\mathcal{F}_{i-1})=_{d}\text{Bin}(n-(i-1)-Y_{i-1},p)\leq_{sd}\text{Bin}(n-(i-1),p).$
@@
$\mathbb{P}(S_{\lfloor BLT^{1/2}\rfloor}>-2L+1)=\mathbb{P}\Big{(}\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}(\eta_{i}-1)>-2L\Big{)}\leq\mathbb{P}\Big{(}\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}(\delta_{i}-1)>-2L\Big{)}.$
@@
where $c=c(B)>0$ is a constant which depends on $B$. Since $Tp=O(A^{-1}n^{-1/3})$, we see that
@@
Therefore, going back to (21) and setting $\ell\coloneqq\lfloor BLT^{1/2}\rfloor(1-np)+(Mp)/2-3L$, we can bound
@@
A simple computation shows that, for all large enough $A$ and $n$, since $LT\geq(1-o(1))n2^{-1}$,
$(Mp)/2-4L\geq\frac{B^{2}L^{2}T}{4n}-4L=L\Big{(}\frac{B^{2}LT}{4n}-4\Big{)}\geq L(B^{2}/12-4),$